OpenStudy (anonymous):

A. 13/35=1/5, 27/73= 2/3, 16/64= 1/4 B. 4/5+2/3= 6/8, 2/5+3/4=5/9, 7/8+1/3=8/11 C. 67/8 - 25/4=10/4, 43/8 - 8/3= 16/5, 16/7- 4/3= 5/4 D. 2/3(3)= 6/9, 1/4(6)= 6/24, 4/5(2)= 8/10

OpenStudy (anonymous):

sorry you have to find the error patterns for each fraction.

OpenStudy (aroub):

Radiance?

OpenStudy (anonymous):

yes

OpenStudy (aroub):

ok, cool :) umm.. I really dont understand your question! I mean what do we suppose to do?

OpenStudy (aroub):

"you have to find the error patterns for each fraction. " oh sorry just saw this!lol sorry! ok i got it :)

OpenStudy (anonymous):

The instructions are to explain the errors in each of the following set. From what I see they were all improperly divided wrong.

OpenStudy (anonymous):

Oh ok no problem

OpenStudy (aroub):

A. 13/35=1/5, 27/73= 2/3, 16/64= 1/4 ^ ^ these two are wrong! okay can you give me a sec, just to solve the rest on a paper! it's better...ok? :)

OpenStudy (anonymous):

No problem

OpenStudy (aroub):

ok just to make sure... have you solved any of these?

OpenStudy (anonymous):

yes I have multiply them by the receiprical and the answers I have gotten dont match with what is written, so therfeore the fractions weren't divided right I suppose

OpenStudy (aroub):

you're talking about B. right?

OpenStudy (anonymous):

for A i mutiplied the 13/35 and got 13/7, so that means 13/35 \[\neq\] 1/5

OpenStudy (anonymous):

same rules apply to the next fraction in that sequence. the only correct fraction is 16/64

OpenStudy (aroub):

yes

OpenStudy (anonymous):

For b you can't just add fractions together without a LCD

OpenStudy (aroub):

but for B, C and D there is NO answers! they are all wrong! yes I'v done it all with LCD

OpenStudy (aroub):

for this one : 4/5+2/3 i got 22/15 and for this one : 2/5+3/4 i got 23/20 and fr this one:7/8+1/3 i got 29/24

OpenStudy (aroub):

and for C. 67/8 - 25/4=17/8, 43/8 - 8/3= 65/24, 16/7- 4/3=20/21

OpenStudy (aroub):

and D. 2/3(3)=2, 1/4(6)= 3/2, 4/5(2)= 8/5

OpenStudy (aroub):

hmm..so what do you think?

OpenStudy (anonymous):

that the answer is the same for all... you cant simplify any of them without having the proper denominator and numerato

OpenStudy (aroub):

what do you mean by "the answer is the same for all" ?

OpenStudy (anonymous):

For b,c,and d the ondividual dindt find the LCD

OpenStudy (aroub):

yes...

OpenStudy (aroub):

well actually ALL!

OpenStudy (anonymous):

for a when you multiply by the reciperacal you get a different denominator so that means the denominator werent changed to get 1/5

OpenStudy (aroub):

hmm... you're right

OpenStudy (aroub):

they just added them or subtracted them without using the LCD...

OpenStudy (anonymous):

ok